I need to simplify these 3 equations?
4+3*5^2
2-4^3 / 8+10
(6^2-2^3)-9 / 3

Answer by Alan3354(69443)

I need to simplify these 3 equations?
4+3*5^2
2-4^3 / 8+10
(6^2-2^3)-9 / 3
-----------------------
They're not equations, they don't have an equal sign.
-------
4+3*5^2
= 4 + 3*25
= 4 + 75
= 79
------------------
2-4^3 / 8+10
This is ambiguous. The DEN might be 8, or 8+10
Add parentheses.
---------------
(6^2-2^3)-9 / 3
= (36 - 8)-9 / 3
Does the 3 divide the 9? Or all the terms?
Use parentheses.